Indicate the equation of the line through (2, -4) and having slope of 3/5.
Answer by stanbon(75887)

Indicate the equation of the line through (2, -4) and having slope of 3/5.
-------
Form:: y = mx + b
Solve for "b"::
-4 = (3/5)*2 + b
----
b = -4 - (6/5)
---
b = -26/5
--------------
Equation:
y = (3/5)x - (26/5)
----------------
Cheers,
Stan H.
------------